# Microphysical Origins of Gravitational Potentials in Astrophysical Systems

### v1

@@ Abstract
Most astrophysical systems (except for very compact objects such as, e.g., black holes and neutron stars) in our universe are characterized by shallow gravitational potentials, with dimensionless compactness $\left|\Phi\right|\equiv r_{s}/R\ll 1$, where $r_{s}$ and $R$ are their Schwarzschild radius and typical size, respectively. While the existence and characteristic scales of such virialized systems depend on gravity, we demonstrate that the value of $\Phi$—and thus the non-relativistic nature of most astrophysical objects—arises from microphysical parameters, specifically the fine structure constant and the electron-to-proton mass ratio, and is fundamentally independent of the gravitational constant $G$. It then follows that large-scale astrophysical velocities and escape velocities associated with virialized systems are determined by electromagnetic and atomic physics, not by gravitation, and that the compactness $\Phi$ is always set by microphysical scales — even for the most compact objects, such as neutron stars, where $\Phi$ is determined by quantities like the pion-to-proton mass ratio. This observation, largely overlooked in the literature, explains why the universe is not dominated by relativistic, compact objects and connects the low entropy of the observable universe to underlying microphysics. Our results emphasize the central but underappreciated role played by dimensionless microphysical constants in shaping the macroscopic gravitational landscape of the universe. In particular, we clarify that this independence of $\Phi$ from $G$ applies specifically to entire, virialized, naturally formed astrophysical systems — such as stars, galaxies, and planets — that have reached equilibrium through self-gravity and microphysical processes. In contrast, arbitrary non-virialized subsystems (e.g., a piece cut from a planet) do not exhibit this property.
@@ I Introduction
The sensitivity of physical processes to the values of fundamental constants, such as the fine-structure constant $\alpha_{e}$ and the proton-to-electron mass ratio $\mu\equiv m_{p}/m_{e}$ — where $m_{p}$ and $m_{e}$ are the proton and electron masses, respectively — has been the subject of extensive theoretical and observational investigation. Variations in these constants impact not only particle physics, but also astrophysical processes, the formation of complex structures, and the potential habitability of the Universe.
@@
These and other explorations of how fundamental constants impact astrophysics and cosmology motivate us to question their possible—if sometimes indirect or hidden—role in shaping the velocity fields associated with typical astrophysical systems. Remarkably, velocities observed in such systems never exceed a threshold of a few thousand km/sec. For example, the Earth’s velocity with respect to the Hubble frame is $\lesssim 400$ km/sec. The escape velocities from Venus, Earth, Mars, Saturn, and the Sun are approximately $6$, $11$, $5$, $35$, and $620$ km/sec, respectively. The solar system velocity relative to the Galactic center is 220220220220240240240240 km/sec, and typical infall velocities of merging clusters and subclusters have never been observed to exceed a few thousand km/sec. The corresponding gravitational wells are characterized by $\Phi+\beta^{2}\sim 0$, where $\Phi$ is the gravitational potential in speed of light-squared, $c^{2}$, units ($\Phi\equiv r_{s}/R$ is the “compactness” parameter where $r_{s}$ and $R$ are its Schwarzschild radius and actual size, respectively), and $\beta$ is the velocity in speed of light units.
@@
This fact is somewhat puzzling, as a much more probable state would be a Universe teeming with black holes (BHs) and neutron stars (NSs), with $|\Phi|\lesssim 1$ and typical velocities approaching the speed of light. Such a Universe would be thermodynamically closer to the final heat death state, in which all BHs are evaporated via Hawking radiation and the far-future asymptotic de Sitter (dS) space is filled with blackbody radiation characterized by the corresponding temperature $T_{dS}\sim 10^{-32}$ eV. Regardless of its initial state, the penultimate state of the Universe is expected to be populated by BHs and NSs. On small scales, such a Universe would be much less homogeneous than the one we actually observe; higher-entropy gravitational systems are simply more clustered. The puzzle, then, is why our Universe is so regular and homogeneous ($|\Phi|\ll 1$). A key insight developed in this work is that, for entire, virialized, and naturally-formed astrophysical systems—such as stars, planets, and galaxies—the dimensionless $\Phi$ is fundamentally set by microphysical parameters (such as $\alpha_{e}$, $\mu$, and other dimensionless ratios), and is independent of the value of the gravitational constant $G$. This property does not apply to arbitrary subsystems (for example, a fragment cut from a planet), but holds for systems that have reached equilibrium through self-gravity and the relevant microphysical processes. Although the dependence of astrophysical masses and radii on fundamental constants is well explored, the general $G$-independence of $\Phi$ for virialized systems is not explicitly discussed in the literature, and we aim to highlight and clarify this point throughout the paper.
@@
This paper is organized as follows. In Section II, we summarize a few basic results from the literature on the typical masses, radii, and temperatures of certain virialized astrophysical systems. In Section III, we highlight the non-observability of the gravitational constant in the absence of a fixed reference scale, and we summarize in Section IV. Throughout, we adopt natural units where the Boltzmann constant $k_{B}$, Planck constant $\hbar$, and the speed of light $c$ are all set to unity. Our estimates are only crude order-of-magnitude calculations, but this is sufficient for our purposes.
@@ II Dimensional Analysis Considerations
Quite generally, the average energy density, $\rho$, and temperature, $T$, characterizing an astrophysical system of mass $M$ and radius $R$, are given by
@@
$GM/R$ $\sim$ $r_{s}/R\sim\Phi\sim T/m_{p}.$ (2)
@@
If both $\rho$ and $T$ on the right-hand sides of Eqs. (1) and (2) are determined by the relevant microphysics of the system in question, then the system of Eqs. (1) and (2) can be solved for $R$, resulting in $R\sim\left(\frac{Tm_{p}}{\alpha_{g}\rho}\right)^{1/2}$ where $\alpha_{g}\equiv Gm_{p}^{2}$. It then follows that $M\sim\rho\left(\frac{Tm_{p}}{\alpha_{g}\rho}\right)^{3/2}$. Crucially, Eq. (2), $\Phi\sim T/m_{p}$, implies that the compactness $\Phi$—not only the density $\rho$—is completely decoupled from $G$. The latter only affects the typical values of $M$ and $R$. For example, assuming $\rho$ and $T$ are fully determined by miscrophysics and are independent of $G$, it immediately follows that $M\propto G^{-3/2}$ and $R\propto G^{-1/2}$. Exceptions are galaxies (and potentially galaxy clusters as well), whose typical $\rho$ does depend on $G$, and so $M$ and $R$ have different $G$-dependencies. Nevertheless, since even in these systems $T$ is independent of $G$, then $\Phi\sim T/m_{p}$ is independent of $G$, and the $G$-independence of the corresponding escape/virial velocities, gravitational lensing and gravitational redshift automatically follows.
@@
The temperature $T$ is a property of the system. For example, the temperature of planets cannot exceed the Rydberg energy (in fact, it must be much lower; see [13]), the temperature of fuel burning stars cannot be smaller than $\sim 1$ keV, otherwise the Coulomb barrier cannot be overcome to allow for nuclear burning, etc. The idea is that temperature of virialized systems sets a lower limit on the escape velocity, and in case that the thermal velocity exceeds the escapse velocity the system quickly disperses and disintegrates. However, the escape velocity sets the gravitational potential, i.e., the ratio $M/R$. If in addition $\rho\sim M/R^{3}$ is known from the microphysics of the system in question, then as argued above, these two constraints can be solved for the characteristic $M$ and $R$.
Clearly, this conclusion is not inconsistent with the fact that $\Phi=GM/R$, and so for arbitrarily chosen $M$ and $R$, the gravitational potential $\Phi$ (and thus escape velocities and gravitational redshift) does depend on $G$. In other words, if we cut a piece of size $R_{0}$ from, e.g., a planet, its energy density does not change, but if $R_{0}$ (and consequently $M\sim\rho R_{0}^{3}$) is arbitrarily chosen, then $\Phi\sim G\rho R_{0}^{2}$ does depend on our arbitrary choice of $R_{0}$ as well as on $G$.
@@
Two immediate consequences of all this are as follows. First, ’line broadening’ $\sigma_{\nu}=\sqrt{\frac{kT}{mc^{2}}}\nu_{0}\sim v\nu_{0}$ around a frequency $\nu_{0}$, caused by thermal motion of the emitters in a gravitationally bound system—essentially a redshift smearing effect—is independent of $G$. Second, although gravitational lensing is sourced by gravity, it is now trivial to see that the deflection angle is independent of $G$. This simply follows from the fact that the geodesic equations depend only on $\Phi$, but the latter is independent of $G$, and it then readily follows that the deflection angle is independent of $G$. $G$ could be a billion times larger or smaller than its measured value, yet the deflection angle would not change. More formally, the deflection angle $\alpha$ is determined by the transverse impulse, i.e., the ratio of the transverse photon momentum change to its momentum, $\alpha=2\int\nabla_{\perp}\Phi\,dl$. Now, since $\Phi$ is typically independent of $G$ in virialized astrophysical systems, then whatever $G$-dependence length scales might have, the combination $\alpha=2\int\nabla_{\perp}\Phi\,dl$ remains independent of $G$, as is appropriate for a geometric effect.
@@
Let us inspect a few examples more closely. Our first example is rocky habitable planets. Following the arguments of [13], the temperature in habitable planets cannot exceed a certain threshold, specifically some fraction $\epsilon<1$ of the Rydberg energy, $Ry\equiv\frac{m_{e}}{2}\alpha_{e}^{2}$, so as not to disrupt the chemistry required for sustaining living organisms, i.e., $|\Phi|\sim v_{esc}^{2}\lesssim T/m_{p}$ where $T=Ry$. All this implies, e.g., [13], that $|\Phi|\sim\epsilon\alpha_{e}^{2}\mu^{-1}$. In other words, since $T<1$ eV and $m_{p}\lesssim 1$ GeV, then $|\Phi|=r_{s}/R\lesssim 10^{-9}$. For planet Earth, with radius $R\lesssim 10^{4}$ km, it implies that the corresponding Schwarzschild radius is $r_{s}\sim 1$ cm. The typical energy density of planets is given by $\rho\sim m_{p}/a_{0}^{3}$, where $a_{0}\equiv(m_{e}\alpha_{e})^{-1}$ is the Bohr radius, i.e., each Bohr volume is populated by approximately a single atom. Together, these simple atomic physics-based estimates of $T$ and $\rho$ uniquely determine $R$ and $M$ as described in Equations (1) and (2). In the case of planets in general, we require $T\sim Ry$ because the atomic structure of matter has to be preserved, albeit the chemistry of living organisms is no longer required to be protected. Thus, we require $T\sim Ry$ and so $\Phi\sim\alpha_{e}^{2}\mu^{-1}$ in this case, and $R$ and $M$ are correspondingly larger than in the habitable planets case by factors of $\epsilon^{-1/2}$ and $\epsilon^{-3/2}$, respectively.
@@
In [15], the estimate of galactic-scale systems required that for a gas cloud to fragment and form stars, the typical temperature has to drop below the Rydberg temperature. This by itself determines $|\Phi|\sim\alpha_{e}^{2}\mu^{-1}$. Another constraint used in [15] was that the dynamical gravitational time is comparable to the (Thomson) cooling time. Together, these constraints resulted in a fair estimate of $M$ and $R$. However, this early analysis ignored the role of CDM. It was revised later in [16] to account for CDM. As expected, the conclusion that $|\Phi|$ is shallow is robust against this improved analysis.
@@
Molecular clouds do not seem to have been discussed in the literature in the context of typical scale estimates. Very general dimensional arguments imply that the vibrational and rotational molecular energy levels are given by $E_{vib}\sim\alpha_{e}^{2}m_{e}\mu^{-1/2}$ and $E_{rot}\sim\alpha_{e}^{2}m_{e}\mu^{-1}$. Considering virialized molecular clouds, the corresponding compactness parameters $\Phi$ are then $|\Phi|\sim\alpha_{e}^{2}\mu^{-3/2}$ and $|\Phi|\sim\alpha_{e}^{2}\mu^{-2}$, respectively, once again demonstrating the shallowness of $|\Phi|$. These estimates are general and decoupled from any assumptions pertaining to the corresponding $M$ and $R$.
@@
White dwarfs (WD) and neutron stars (NS) are held against their own gravity by degeneracy pressure. In the case of WD, the degeneracy is that of electrons with typical spacing of order the Compton wavelength of electrons, naively $T\sim m_{e}$, and the corresponding $\Phi\sim\mu^{-1}$. In the case of NS, the effective temperature is $m_{\pi}$ and so $\Phi\sim m_{\pi}/m_{p}\lesssim 1$. Thus, only when gravity is opposed by the strong interaction, or by no opposing force at all, is $|\Phi|\sim 1$ and the corresponding escape velocities become relativistic.
@@
These results, along with others based on references [13, 14, 15, 16] for various virialized astrophysical systems, are compiled and summarized in Table 1. The table presents typical masses, radii, number densities, and compactness in terms of the fine-structure constant $\alpha_{e}$, the proton-to-electron mass ratio $\eta\equiv m_{p}/m_{e}$, the proton-to-pion mass ratio $\kappa\equiv m_{p}/m_{\pi}$, the relative electromagnetic-to-gravitational interaction strength $\lambda\equiv\alpha_{e}/\alpha_{g}$ (where $\alpha_{g}\equiv Gm_{p}^{2}$), and a dimensionless parameter $\epsilon\sim 0.1$.
@@
It is worth noting that, while the primary focus of the present work is the rightmost column showing the compactness parameter for a few typical astrophysical systems, the information in the other columns illustrates the valuable insights gained from basic dimensional analysis. For example, planets are less massive than their host stars fundamentally because $\alpha_{e}<1$, NS’s are less massive than their progenitor stars because $m_{\pi}<m_{p}$, and the observed Sun-to-Jupiter mass ratio, $O(10^{3})$, is simply determined by $\alpha_{e}^{-3/2}$. Likewise, the star-to-habitable planet (such as Earth) mass ratio is given by $\sim(\epsilon\alpha_{e})^{-3/2}$. While these observed ratios are typically explained by the details of various astrophysical processes, these processes are ultimately governed by dimensionless ratios of fundamental constants of nature, as demonstrated here.

### v2

# Extensive and Intensive Aspects of Astrophysical Systems and Fine-Tuning

@@ Abstract
Most astrophysical systems (except for very compact objects such as, e.g., black holes and neutron stars) in our Universe are characterized by shallow gravitational potentials, with dimensionless compactness $|\Phi|\equiv r_{s}/R\ll 1$, where $r_{s}$ and $R$ are their Schwarzschild radius and typical size, respectively. While the existence and characteristic scales of such virialized systems depend on gravity, we demonstrate that the value of $|\Phi|$ – and thus the non-relativistic nature of most astrophysical objects – arises from microphysical parameters, specifically the fine structure constant and the electron-to-proton mass ratio, and is fundamentally independent of the gravitational constant $G$. In fact, the (generally extensive) gravitational potential becomes ‘locally’ intensive at the system boundary; the compactness parameter corresponds to the binding energy (or degeneracy energy, in the case of quantum degeneracy pressure-supported systems) per proton, representing the amount of work that needs to be done in order to allow proton extraction from the system. More generally, extensive properties of gravitating systems depend on $G$, whereas intensive properties do not. It then follows that peak rms values of large-scale astrophysical velocities and escape velocities associated with naturally formed astrophysical systems are determined by electromagnetic and atomic physics, not by gravitation, and that the compactness $|\Phi|$ is always set by microphysical scales – even for the most compact objects, such as neutron stars, where $|\Phi|$ is determined by quantities like the pion-to-proton mass ratio. This observation, largely overlooked in the literature, explains why the Universe is not dominated by relativistic, compact objects and connects the relatively low entropy of the observable Universe to underlying basic microphysics. Our results emphasize the central but underappreciated role played by dimensionless microphysical constants in shaping the macroscopic gravitational landscape of the Universe. In particular, we clarify that this independence of the compactness $|\Phi|$ from $G$ applies specifically to entire, virialized or degeneracy pressure-supported systems, naturally formed astrophysical systems – such as stars, galaxies, and planets – that have reached equilibrium between self-gravity and microphysical processes. In contrast, arbitrary subsystems (e.g., a piece cut from a planet) do not exhibit this property; well within/outside the gravitating object the rms velocity is suppressed and $G$ reappears. Finally, we point out that a clear distinction between intensive and extensive astrophysical/cosmological properties could potentially shed a new light on the mass hierarchy and the cosmological constant problems; both may be related to the large complexity of our Universe.
@@ I Introduction
The sensitivity of physical processes to the values of fundamental constants, such as the fine-structure constant $\alpha_{e}$ and the proton-to-electron mass ratio $\mu\equiv m_{p}/m_{e}$ – where $m_{p}$ and $m_{e}$ are the proton and electron masses, respectively – has been the subject of extensive theoretical and observational investigation. Variations in these constants impact not only particle physics, but also astrophysical processes, the formation of complex structures, and the potential habitability of the Universe.
@@
These and other explorations of how fundamental constants impact astrophysics and cosmology motivate us to question their possible – if sometimes indirect or hidden – role in shaping the velocity fields associated with typical astrophysical systems. Remarkably, velocities observed in such systems never exceed a threshold of a few thousand km/sec. For example, the Earth’s velocity with respect to the Hubble frame is 400$\lesssim$ 400≲ 400 km/sec. The escape velocities from Venus, Earth, Mars, Saturn, and the Sun are approximately 666, 111111, 555, 353535, and 620620620 km/sec, respectively. The solar system velocity relative to the Galactic center is 220220220240240240 km/sec, and typical infall velocities of merging clusters and subclusters have never been observed to exceed a few thousand km/sec. The corresponding gravitational wells are characterized by $\Phi+\beta^{2}\sim 0$, where $\Phi$ is the gravitational potential in speed of light-squared, $c^{2}$, units ($|\Phi|\equiv r_{s}/R$ is the “compactness” parameter where $r_{s}$ and $R$ are its Schwarzschild radius and actual size, respectively), and $\beta$ is the velocity in speed of light units.
@@
This fact is somewhat puzzling, as a much more probable state would be a Universe teeming with black holes (BHs) and neutron stars (NSs), with $|\Phi|\lesssim 1$ and typical velocities approaching the speed of light. This is so simply because the next-to-highest entropy state of gravitating system is when it forms a BH, that eventually evaporates by emitting maximum entropy black body radiation. Such a Universe would be thermodynamically closer to the final heat death state, in which all BHs are evaporated via Hawking radiation and the far-future asymptotic de Sitter dS space is filled with blackbody radiation characterized by the corresponding temperature $T_{dS}\sim 10^{-32}$ eV [11]. Regardless of its initial state, the penultimate state of the Universe is expected to be populated by BH’s and NS’s. On small scales, such a Universe would be much less homogeneous than the one we actually observe; higher-entropy gravitational systems are simply more clustered. The puzzle, then, is why our Universe is so regular and homogeneous ($|\Phi|\ll 1$). A key insight developed in this work is that, for entire, virialized (or supported by degeneracy pressure), and naturally-formed astrophysical systems – such as stars, planets, and galaxies – the dimensionless $|\Phi|$ is fundamentally set by microphysical parameters (such as $\alpha_{e}$, $\mu$, and other dimensionless ratios), and is independent of the value of the gravitational constant $G$; given the fact that $\alpha_{e}\ll 1$ and $\mu\gg 1$ the Universe is simply not sufficiently old for having all the potentially available nuclear fuel exhausted by now to allow for stars to collapse, the atomic scale that determine the mass density of rocky planets is much larger than nuclear scale, etc. In other words, non-gravitational interactions are sufficiently effective in inhibiting the Universe from reaching its ultimate fate even 14$\sim$ 14∼ 14 Gyrs after the big bang. The intensive nature of $|\Phi|$ does not apply to arbitrary subsystems (for example, a fragment cut from a planet), but holds for systems that have reached equilibrium through self-gravity and the relevant microphysical processes. Although the dependence of astrophysical masses and radii on fundamental constants is well explored, the general $G$-independence of $|\Phi|$ for virialized or degeneracy-pressured supported systems is not explicitly discussed in the literature, and we aim to highlight and clarify this point throughout the paper.
On a different front, and seemingly a sharp digression, it is worth noting that viewing the Planck mass $m_{pl}$ and gravitational masses $m_{gr}$ through the lens of extensive and intensive properties – contrasted with the Higgs mass $m_{H}$ and inertial masses $m_{in}$ – could potentially shed new light on the longstanding mass hierarchy problem, e.g. [12, 13, 14, 15], spanning $\sim$17 orders of magnitude, as well as the cosmological constant problem, e.g [16, 17, 18, 19], of $\sim$122 orders of magnitude. From this perspective, $m_{pl}^{3}$ and $m_{gr}^{3}$ may be regarded as extensive quantities, while the ratio $m_{gr}/m_{pl}$ remains intensive. Such a perspective leaves the standard gravitational framework intact while potentially offering fresh insights into why these fine-tuning ’problems’ arise in the first place.
@@
This paper is organized as follows. In Section II, we summarize a few basic results from the literature on the typical masses, radii, and temperatures of certain virialized or degeneracy-pressure supported astrophysical systems. In Section III we discuss the relation between extensive formulation of gravity and local scale invariance and how the latter emerges when intensive microscopic description is generalized to extensive formulation of gravity. In Section IV we discuss possible implications of extensive Planck mass and we summarize in Section V. Throughout, we adopt natural units where the Boltzmann constant $k_{B}$, Planck constant $\hbar$, and the speed of light $c$ are all set to unity. Our estimates are only crude order-of-magnitude calculations, but this is sufficient for our purposes.
@@ II Dimensional Analysis Considerations
In the following we start with the basic equations to correctly put the context and for future reference but quickly move on to dimensional considerations. The classical equations for gravitating systems include the Poisson equation for the gravitational potential $\Phi$,
$\nabla^{2}\Phi=4\pi G\rho,$ (1)
and the continuity and Euler equations, respectively
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\mathbf{v})=0,$ (2)
$\frac{\partial\mathbf{v}}{\partial t}+(\mathbf{v}\cdot\nabla)\mathbf{v}=-\frac{\nabla P}{\rho}-\nabla\Phi,$ (3)
supplemented by the equation of state (EOS)
$P=P(\rho,T,\{\chi_{i}\}).$ (4)
The latter relates the pressure, $P$, to the mass density, $\rho$, and possibly other parameters denoted by $\{\chi_{i}\}$. Here $\mathbf{v}$ is the velocity field. In realistic systems $P$ is provided primarily by non-gravitational interactions or otherwise by quantum degeneracy. However, at the simple-minded level of discussion adopted here we model the astrophysical system as purely gravitational although the mass density values we adopt (as will be seen in the following) certainly rely on the existence of non-gravitational interactions. To support these over-idealized spherically-symmetric, constant density and static systems against their own gravitational pull without recourse to non-gravitationally produced pressure we are automatically lead to model our spherically symmetric static systems as de-Sitter solutions of the Einstein field equations in Section III.2.
In the following, $E_{\star}$ is either the (electromagnetic) binding energy per atom, e.g. in the case of planets, nuclear burning stars, galaxies, etc., or the quantum degeneracy pressure in the case of WD’s or NS’s. The mass density and compactness of astrophysical system of mass $M$ and radius $R$, are given by
@@
$GM/R$$\sim$ $r_{s}/R\sim|\Phi|\sim E_{\star}/m_{p},$ (6)
where the electromagnetic binding energy is consistent with the virial temperature of the system, $E_{\star}\sim T$, and $r_{s}\equiv 2GM$ is the Schwarzschild radius (strictly, it would be the corresponding de Sitter scale in our approximation but given that the current analysis is only an order of magnitude approximation the distinction is of no significant relevance). It should be stressed that whereas $\rho$ is a local property of the astrophysical system, the gravitational potential is not, and $|\Phi|\sim E_{\star}/m_{p}$ expresses values at the boundaries of entire, naturally-formed, systems, with typical size $R$. For such systems $E_{\star}/m_{p}\sim|\Phi|\sim G\rho R^{2}$ it follows that $R\propto G^{-\frac{1}{2}}$ provided that $\rho$ is indeed determined exclusively by microphysics. In this case $M\propto G^{-\frac{3}{2}}$, and so does $N\propto G^{-\frac{3}{2}}$ – the number of protons which compose the system (assuming negligible binding energy). However, cutting a piece or fragment of size $\tilde{R}$ from the system, e.g. a rock from a planet, then whereas $\rho$ is unchanged $|\Phi(\tilde{R})|\sim G\rho\tilde{R}^{2}<G\rho R^{2}\sim|\Phi(R)|$ is clearly not G-independent for arbitrarily chosen $\tilde{R}$. It addition, the exterior solution for $r>R$ is $|\Phi|\sim GM/r=\Phi R/r$, and so the maximum value of the gravitational potential at $r=R$ is G-independent. The unique ‘status’ of the system boundary that allows for the decoupling of $G$ from $|\Phi|$ at this surface is physically clear; typical velocities at the system boundary must not exceed the escape velocity. The latter is bounded by the thermal velocity which is in turn set by microphysics.
@@
In case that the compactness is determined by the binding energy, then since $|\Phi|=v_{esc}^{2}$, the requirement that the escape velocity $v_{esc}$ is larger than the thermal velocity $v_{th}$, and that the rms of the latter is determined by the temperature and the mass of the heaviest stable particle, the proton, via $m_{p}v_{th}^{2}\sim T$, we obtain that typically $|\Phi|\sim T/m_{p}$. If both $\rho$ and $E_{\star}$ on the right-hand sides of Equations 5 and 6 are determined by the relevant microphysics of the system in question, then the system of Equations 5 and 6 can be solved for $R$, resulting in $R\sim\left(\frac{E_{\star}m_{p}}{\alpha_{g}\rho}\right)^{1/2}$ and $\alpha_{g}\equiv Gm_{p}^{2}$. It then follows that $M\sim\rho\left(\frac{E_{\star}m_{p}}{\alpha_{g}\rho}\right)^{3/2}$. Crucially, Equation 6, $|\Phi|\sim E_{\star}/m_{p}$, implies that the compactness $|\Phi|$ – not only the density $\rho$ – is completely decoupled from $G$. The latter only affects the typical values of $M$ and $R$. As pointed out above, assuming $\rho$ and $E_{\star}$ are fully determined by microphysics and are independent of $G$, it immediately follows that $M\propto G^{-3/2}$ and $R\propto G^{-1/2}$. Exceptions are galaxies (and potentially galaxy clusters as well), whose typical $\rho$ does depend on $G$, and so $M$ and $R$ have different $G$-dependencies. Nevertheless, since even in these systems $T$ is independent of $G$, then $|\Phi|\sim T/m_{p}$ is independent of $G$, and the $G$-independence of the corresponding escape/virial velocities, maximum level of gravitational lensing and gravitational redshift automatically follows.
@@
The temperature $T$ is a property of the system. For example, the temperature of planets cannot exceed the Rydberg energy (in fact, it must be much lower; see [22]), the temperature of fuel burning stars cannot be smaller than 1$\sim$ 1∼ 1 keV, otherwise the Coulomb barrier cannot be overcome to allow for nuclear burning, etc. The idea is that temperature of virialized systems sets a lower limit on the escape velocity, and in case that the thermal velocity exceeds the escape velocity the system quickly disperses and disintegrates. However, the escape velocity sets the gravitational potential, i.e., the ratio $M/R$. If in addition $\rho\sim M/R^{3}$ is known from the microphysics of the system in question, then as argued above, these two constraints can be solved for the characteristic $M$ and $R$. It is perhaps worth mentioning that since $|\Phi|\sim E_{\star}/m_{p}$, then the highest possible temperatures of naturally formed astrophysical systems cannot exceed $10^{13}$, otherwise $|\Phi|$ reaches unity, the escape velocity is close to the speed of light, and the system quickly disperses. For comparison, typical temperatures of the intracluster plasma in galaxy clusters fall in the range $\sim 10^{7}-10^{8}$, accretion disks around BH’s and NS’s are usually not hotter than $10^{9}$, the temperatures in cores of newly formed NS’s are typically $10^{11}$, etc.
@@
Two immediate consequences of all this are as follows. First, ’line broadening’ $\sigma_{\nu}=\sqrt{\frac{kT}{mc^{2}}}\nu_{0}\sim v\nu_{0}$ around a frequency $\nu_{0}$, caused by thermal motion of the emitters in a gravitationally bound system – essentially a redshift smearing effect – is maximized at the boundaries of astrophysical systems in a G-independent fashion. Second, although gravitational lensing is sourced by gravity, it is now trivial to see that the maximum deflection angle is independent of $G$. This simply follows from the fact that the geodesic equations depend only on $|\Phi|$, but the maximum absolute value of the latter is independent of $G$, and it then readily follows that the maximal deflection angle is independent of $G$. The latter could be a billion times larger or smaller than its measured value, yet the maximum value of the deflection angle would not change for it is set by $G$-independent microphysics. More formally, the deflection angle $\alpha$ is determined by the transverse impulse, i.e., the ratio of the transverse photon momentum change to its momentum, $\alpha=2\int\nabla_{\perp}\Phi\,dl$. Now, since $\Phi$ is typically independent of $G$ in virialized and degeneracy-pressure supported astrophysical systems, then whatever $G$-dependence length scales might have, the combination $\alpha=2\int\nabla_{\perp}\Phi dl$ remains independent of $G$ at its maximum.
@@
In the following we inspect a few examples more closely. Our first example is rocky habitable planets. Following the arguments of [22], the temperature in habitable planets cannot exceed a certain threshold, specifically some fraction $\epsilon<1$ of the Rydberg energy, $Ry\equiv\frac{m_{e}}{2}\alpha_{e}^{2}$, so as not to disrupt the chemistry required for sustaining living organisms, i.e., $|\Phi|\sim v_{esc}^{2}\lesssim T/m_{p}$ where $T=Ry$. All this implies, e.g., [22], that $|\Phi|\sim\epsilon\alpha_{e}^{2}\mu^{-1}$. In other words, since $T<1$ eV and $m_{p}\lesssim 1$ GeV, then $|\Phi|=r_{s}/R\lesssim 10^{-9}$. For planet Earth, with radius $R\lesssim 10^{4}$ km, it implies that the corresponding Schwarzschild radius is $r_{s}\sim 1$ cm. The typical energy density of planets is given by $\rho\sim m_{p}/a_{0}^{3}$, where $a_{0}\equiv(m_{e}\alpha_{e})^{-1}$ is the Bohr radius, i.e., each Bohr volume is populated by approximately a single atom. Together, these simple atomic physics-based estimates of $T$ and $\rho$ uniquely determine $R$ and $M$ as described in Equations 5 and 6. In the case of planets in general, we require $T\sim Ry$ because the atomic structure of matter has to be preserved, albeit the chemistry of living organisms is no longer a condition that needs to be satisfied. Thus, we require $T\sim Ry$ and so $|\Phi|\sim\alpha_{e}^{2}\mu^{-1}$ in this case, and $R$ and $M$ are correspondingly larger than in the habitable planets case by factors of $\epsilon^{-1/2}$ and $\epsilon^{-3/2}$, respectively.
@@
In [24], the estimate of galactic-scale systems required that for a gas cloud to fragment and form stars, the typical temperature has to drop below the Rydberg temperature. This by itself determines $|\Phi|\sim\alpha_{e}^{2}\mu^{-1}$. Another constraint used in [24] was that the dynamical gravitational time is comparable to the (Thomson) cooling time. Together, these constraints resulted in a fair estimate of $M$ and $R$. However, this early analysis ignored the role of cold dark matter (CDM). It was revised later in [25] to account for CDM. As expected, the conclusion that $|\Phi|$ is shallow is robust against this improved analysis.
@@
Molecular clouds do not seem to have been discussed in the literature in the context of typical dimensional analysis-based scale estimates. Very general dimensional arguments imply that the vibrational and rotational molecular energy levels are given by $E_{vib}\sim\alpha_{e}^{2}m_{e}\mu^{-1/2}$ and $E_{rot}\sim\alpha_{e}^{2}m_{e}\mu^{-1}$. Considering virialized molecular clouds, the corresponding compactness parameters $|\Phi|$ are then $|\Phi|\sim\alpha_{e}^{2}\mu^{-3/2}$ and $|\Phi|\sim\alpha_{e}^{2}\mu^{-2}$, respectively, once again demonstrating the shallowness of $|\Phi|$. These estimates are general and decoupled from any assumptions pertaining to the corresponding $M$ and $R$.
@@
White dwarfs (WD) and neutron stars (NS) are held against their own gravity by degeneracy pressure. In the case of WD, the degeneracy is that of electrons with typical spacing of order the Compton wavelength of electrons, naively $T\sim m_{e}$, and the corresponding $|\Phi|\sim\mu^{-1}$. In the case of NS, the effective temperature is $m_{\pi}$, where $m_{\pi}$ is the pion mass, and so $|\Phi|\sim\kappa^{-1}\lesssim 1$, where $\kappa\equiv m_{p}/m_{\pi}$ is the proton-to-pion mass ratio. Thus, only when gravity is opposed by the strong interaction, or by no opposing force at all, is $|\Phi|\sim 1$ and the corresponding escape velocities become relativistic.
@@
These results, along with others based on references [22, 23, 24, 25] for various virialized and degeneracy-pressure supported astrophysical systems, are compiled and summarized in Table 1. The table presents typical masses, radii, number densities, and compactness in terms of the fine-structure constant $\alpha_{e}$, the proton-to-electron mass ratio $\eta\equiv m_{p}/m_{e}$, the proton-to-pion mass ratio $\kappa\equiv m_{p}/m_{\pi}$, the relative electromagnetic-to-gravitational interaction strength $\lambda\equiv\alpha_{e}/\alpha_{g}$ (where $\alpha_{g}\equiv Gm_{p}^{2}$), and a dimensionless parameter $\epsilon\sim 0.1$.
@@
It is worth noting that, while the primary focus of the present work is the rightmost column showing the compactness parameter for a few typical astrophysical systems, the information in the other columns illustrates the valuable insights gained from basic dimensional analysis. For example, planets are less massive than their host stars fundamentally because $\alpha_{e}<1$, NS’s are less massive than their progenitor stars because $m_{\pi}<m_{p}$, and the observed Sun-to-Jupiter mass ratio, $O(10^{3})$, is simply determined by $\alpha_{e}^{-3/2}$. Likewise, the star-to-habitable planet (such as Earth) mass ratio is given by $\sim(\epsilon\alpha_{e})^{-3/2}$. While these observed ratios are typically explained by the details of various astrophysical processes, these processes are ultimately governed by dimensionless ratios of fundamental constants of nature, as demonstrated here. It is also important to realize that the fact that $|\Phi|\ll 1$ is not a result of the relative weakness of gravity in comparison to the other interactions gauged by, e.g. $\lambda\equiv\alpha_{e}/\alpha_{g}$, but rather a result of the weakness of the latter in absolute terms, i.e. the fact that $\alpha_{e}\ll 1$, $\mu\gg 1$, etc.